# Notions of entropy for unitary representations

### v2

@@ 8.3 The strong quotient topology on representations
If separable representations $\pi$ and $\rho$ satisfy $\overline{\Sigma_{\bullet}(\pi)}=\overline{\Sigma_{\bullet}(\rho)}$, then they are approximately equivalent. This is a classic weakening of equivalence for representations, and is the reason I propose the term ‘approximate association’ in Subsection 8.1. From penetrating works by Glimm [Gli60] and Voiculescu [Voi76], approximate equivalence is known to have several different characterizations, many of them appearing much stronger than the definition above; see also [Arv77] or [Con00, Section 41]. In those works, our definition above is often called ‘weak approximate equivalence’, and is then proved to imply the other characterizations. However, we do not need this theory below. Approximate equivalence is also the same as ‘weak equivalence in the sense of Zimmer’: see Remark 8.3. The term ‘summary’ is new.
For any topological space $X$, we endow $\mathcal{K}(X)$ with the classic Vietoris topology: see [Eng89, Exercises 2.7.20, 3.12.27 and 4.5.23] or [Kec95, Subsection I.4.F]. We assume standard properties from those references in the sequel. In particular, if $X$ is compact and metrizable, then so is $\mathcal{K}(X)$, for instance using a Hausdorff metric. As a result, in this case we can describe the Vietoris topology in terms of convergence of sequences. More specifically, for any sequence $(K_{n})_{n\geq 1}$ in $\mathcal{K}(X)$, define its topological upper and lower limits to be the sets
@@
respectively. If $X$ is compact and metrizable, then the sequence $(K_{n})_{n\geq 1}$ Vietoris converges if and only if its topological upper and lower limits are equal, and in this case that common set is the Vietoris limit of the sequence.
In our setting, we endow the product space on the right-hand side of (8.5) with the product of the Vietoris topologies. Since each $\Sigma_{k}(\mathfrak{A})$ is compact and metrizable, and the product is countable, this whole product space is again compact and metrizable.
@@
As far as I know, this topology was first studied by Abért and Elek, prompted by an analogous idea in graph theory from [BR11] (see also Remark 8.16 below). Their paper [AE11] focuses on the analogous definition in ergodic theory, but indicates the story for unitary representations of groups as well. However, the term ‘strong quotient’ is new. We comment on this further after presenting some basic properties.
One key conclusion of the representation theoretic analog of [AE11, Theorem 1] is that, if one allows all separable representations, then the image of the summary map is a compact subset of the right-hand side of (8.5). In particular, this image set is complete for any compatible metric on that space. The proof adapts with only cosmetic changes to general unital C*-algebras, and the result has the following consequence.
@@
By (8.1), if $\varphi\in\Sigma_{k}(\mathfrak{A})$, then it is asymptotically associated to $(\pi_{n})_{n\geq 1}$ if and only if it lies in $\mathrm{T}\limsup_{n}\overline{\Sigma_{k}(\pi_{n})}$. The choice of Tlim $\mathrm{T}\limsup$ rather Tlim $\mathrm{T}\liminf$ here is a matter of convention. If $(\pi_{n})_{n\geq 1}$ strong-quotient converges then this choice makes no difference, but otherwise it can be consequential. The next lemma illustrates this.
@@
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. Assume that $\boldsymbol{\pi}$ strong-quotient converges. If $\varphi$ and $\psi$ are disjoint and they are both asymptotically associated to $\boldsymbol{\pi}$, then so is $\mathrm{diag}(\varphi,\psi)$, and so is $\varphi+\psi$ if $k=\ell=1$.
@@
The sum $\varphi+\psi$ is handled in the same way, using Corollary 8.11(a) in place of Lemma 8.9. ∎
Without assuming strong-quotient convergence, the two factors in the Cartesian product above might be nonempty only along disjoint subsequences of indices $n$.
Lemma 8.15.
Proof.
The term ‘strong quotient’ is chosen to locate this topology relative to some other standard notions for representations. Let us quickly review some of these in order to explain this choice.
The most common topology on the spectrum $\widehat{\mathfrak{A}}$ is the Fell topology, which is the pullback of the Jacobson topology on $\mathrm{Prim}\,\mathfrak{A}$. This has several equivalent definitions, originally explored by Godemont [God48] and Fell [Fel60b]: see [Dix77, Chapter 3] or [Fol16, Section 7.2].
One can try to extend this topology to larger spaces of representations than just irreducible ones. But then the different definitions no longer coincide in general. Two possibilities were defined and studied by Fell in [Fel60a, Section 3] and [Fel62]. He called them the ‘quotient’ and ‘inner hull-kernel’ topologies. Modern practice usually assigns Fell’s name to his ‘inner hull-kernel topology’, but it is actually the quotient topology that is closer to our present needs.
These topologies can be defined in terms of the sets of completely positive maps associated to representations. The difference boils down to whether one allows finite sums or convex combinations (giving the Fell topology) or not (giving the quotient topology). This sits alongside our discussion of the difference between weak and approximate association in Subsection 8.4.
Focusing on the quotient topology, it can be defined by pulling back the product of a sequence of topologies on the spaces $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ under the summary map, much like our approach to the strong-quotient topology. But the quotient topology uses the ‘lower Vietoris’ topologies rather than the Vietoris topologies on those spaces (see [Mic51, Appendix]). The lower Vietoris topologies are much weaker: indeed, they are usually not even T1, and so limits of sequences need not be unique. More concretely, a sequence $(\pi_{n})_{n\geq 1}$ converges to another representation $\pi$ in the quotient topology if
@@
So, in this topology, any representation whose $k$-summaries are sufficiently small for every $k$ is a limit of $(\pi_{n})_{n\geq 1}$.
For representations, one manifestation of this weakness is the following. For a set $R$ of representations (or equivalence classes of them) and an element $a\in\mathfrak{A}$, the norm map
@@
is lower semicontinuous but not necessarily continuous for the quotient topology on $R$. See, for instance, the longer discussion in [Fel60b, Chapter II] or [Dix77, Section 3.3] about the case $R=\widehat{\mathfrak{A}}$.
In a different direction, recent research has seen a swell of interest in ‘strong convergence’ of representations. This is defined precisely by the requirement that those norm maps in (8.7) converge to their values on a limiting representation: see, for example, the survey [Mag25].
@@
  • For quotient-topology convergence: for every open subset $O$ of $\Sigma_{k}(\mathfrak{A})$ that meets $\overline{\Sigma_{k}(\pi)}$, and for all sufficiently large $n$, there are some $k$-tuples $V_{n}$ in $H_{\pi_{n}}$ such that $\Phi^{\pi_{n}}_{V_{n}}\in O$.
  • For strong convergence: for every $a\in\mathfrak{A}$, every $\varepsilon>0$, and for all sufficiently large $n$, all unit vectors $v\in H_{\pi_{n}}$ satisfy $\|\pi_{n}(a)v\|\leq\|\pi(a)\|+\varepsilon$.
On the other hand, control of the norm maps in (8.7) does not require controlling multiplicities, or tuples of more than one vector. In fact, the strong convergence of an AP sequence $(\pi_{n})_{n\geq 1}$ is equivalent to Vietoris convergence of the sequence of closed convex hulls $\overline{\mathrm{conv}}\,\Sigma_{1}(\pi_{n})$.
@@
8.4 Almost periodic sequences and asymptotic association
Fix a separable C*-algebra $\mathfrak{A}$. In the next section, the almost periodic entropy of a map $\varphi$ is defined in terms of the volumes of the $O$-typical sets as $O$ ranges over neighbourhoods of $\varphi$. The representations that one uses belong to the following sequences.
@@ 8.4 Almost periodic sequences and asymptotic association
Later we also need another mode of convergence for an AP sequence $\boldsymbol{\pi}$. This is the convergence of the normalized tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ to some limit tracial state $\tau$ in the weak topology of $\mathfrak{A}_{+}^{\ast}$. In the terminology of free probability theory, this asserts that, for any finite subset $F$ of $\mathfrak{A}$, the tuples $(\pi_{n}(a):\ a\in F)$ form a sequence of ‘microstates’ for $F$ in the ‘non-commutative probability space’ $(\mathfrak{A},\tau)$ (see [Voi02], for example). This convergence of tracial states is one of the hypotheses of Theorem C. It determines which tracial state should be used to define the Fuglede–Kadison determinant in the conclusion of that theorem.
@@
After identifying $\overline{\mathrm{tr}}_{kd_{n}}$ with $\overline{\mathrm{tr}}_{d_{n}}\otimes\overline{\mathrm{tr}}_{k}$, this follows from the continuity of the operation $(\cdot)\otimes\overline{\mathrm{tr}}_{k}$ on tracial states. It can be checked entry-wise for an element of $\mathrm{M}_{k}(\mathfrak{A})$. ∎
@@ 9 Almost periodic entropy
In this section we define the almost periodic entropy of an $\mathrm{M}_{k}$-valued completely positive map on a separable C*-algebra. We then build up its properties towards the proof of Theorem C, and follow that with a few consequences.
@@ 9.1 Preliminary results from high-dimensional probability
Then $v(d)$ is equal to the volume of the unit ball in $\mathbf{C}^{d}$ [Rud08, Subsection 1.4.9]. By Stirling’s approximation, this function satisfies
@@
A key feature of the spherical measures $\sigma_{d-1}$ in high dimensions is the phenomenon of measure concentration. We need the following special case.
@@
If $f(v)=\langle Av,v\rangle$ for $v\in\mathrm{S}^{2d-1}$, then $f$ is $2\|A\|$-Lipschitz, and its spherical average is $\overline{\mathrm{tr}}_{d}A$ by the invariance of trace under conjugation. So the desired inequality follows from the concentration of general Lipschitz functions on high-dimensional spheres: see, for instance [MS86, Corollary V.2]
@@
See Lemma 9.13 below for a generalization of Corollary 9.2 that depends on Lemma 9.1 in the same way.
Subsection 1.2 discusses the analogy between the Gram matrix of a tuple of vectors and the joint distribution of a tuple of finite-valued random variables. The next theorem adds another layer to this analogy: a ‘method of types’ interpretation for $\log\det Q$ when $Q$ is a positive semi-definite matrix. It approximates the volumes of the sets $T(d,O)$ when $O$ is a small neighbourhood of $Q$. In information theory, the method of types captures the basic combinatorial meaning of discrete Shannon entropy [CT06, Section 11.1]. The differential entropy of a jointly Gaussian random vector also has a method-of-types interpretation in terms of volumes [CT06, Section 8.2].
@@
    (Lower bound) If $Q$ is nonsingular and $O$ is any neighbourhood of $Q$ in $\mathrm{M}_{k+}$, then
@@
Identifying $\mathrm{M}_{d,k}$ with $\mathbf{C}^{dk}$ and integrating in polar coordinates [Rud08, Subsection 1.4.3], we obtain
@@
By Corollary 9.2 and (9.2), this is greater than $v(d)^{k}$ for all sufficiently large $d$.
@@
If we regard an $d$-by-$k$ matrix as the $d$-tuple of its rows, then right-multiplication by $R$ on $\mathrm{M}_{d,k}$ becomes the direct sum of $d$ copies of $R^{\mathrm{T}}$ acting on $\mathbf{C}^{\oplus k}$. Regarded as a real linear transformation acting on $2kd$ real linear dimensions, this has Jacobian $|\det R|^{2d}$: see, for instance, [Rud08, Subsection 1.3.5]. Therefore (9.5) gives
@@
Since $O^{\prime}$ is an arbitrary neighbourhood of $Q^{\prime}$, this shows that part (a) for $Q^{\prime}$ follows from part (a) for $Q$. Similarly, if $R$ is invertible, then we may reverse the roles of $Q$ and $Q^{\prime}$ above and deduce that part (b) for $Q^{\prime}$ follows from part (b) for $Q$.
@@
$O:=\{Q^{\prime}\in\mathrm{M}_{k+}:\ \langle e_{1},Q^{\prime}e_{1}\rangle<\varepsilon^{2}\ \hbox{and}\ \max_{i}q^{\prime}_{ii}<r^{2}\}.$
Then $O$ is a neighbourhood of $Q$ such that
$T(d,O)=\big{\{}X=[x_{1},\dots,x_{k}]\in\mathrm{M}_{d,k}:\ \|x_{1}\|<\varepsilon\ \hbox{and}\ \max_{i}\|x_{i}\|<r\big{\}},$
and hence
@@
Theorem 9.3 is a template for Theorem C. (It is also a special case of that theorem with $\mathfrak{A}=\mathrm{M}_{k}$.) The rather soft proof given above, which avoids evaluating any integrals over spaces of matrices exactly, is also a precursor to the proof of Theorem C.
For discrete Shannon entropy, the usual proofs in the method of types involve counting strings that have exactly a given empirical distribution and then applying Stirling’s approximation. By contrast, in Step 3 of the proof of Theorem 9.3 we use a change of variables to transport the desired estimate from $I_{k}$ to most other elements of $\mathrm{M}_{k+}$. This way of using the symmetries of $\mathrm{vol}_{2kn}$ has no obvious analog for probability distributions over finite sets. A related use of symmetry is also essential to our later proof of Theorem C.
Further references.
[…]

### v3

@@ 8.3 The strong-quotient topology on representations
Two separable representations $\pi$ and $\rho$ are approximately equivalent if they satisfy $\overline{\Sigma_{\bullet}(\pi)}=\overline{\Sigma_{\bullet}(\rho)}$. This is a classic weakening of equivalence for representations. It is the reason I propose the term ‘approximate association’ in Subsection 8.1. From penetrating works by Glimm [72] and Voiculescu [153], approximate equivalence is known to have several different characterizations, many of them appearing much stronger than the definition above; see also [8] or [38, Section 41]. In those works, the definition we take above is often called ‘weak approximate equivalence’, and it is then proved to imply the other characterizations. However, we do not need this theory below. Approximate equivalence is also the same as ‘weak equivalence in the sense of Zimmer’: see Remark 8.3. Our use of the term ‘summary’ is new.
For any topological space $X$, we endow $\mathcal{K}(X)$ with the classic Vietoris topology: see [54, Exercises 2.7.20, 3.12.27 and 4.5.23] or [92, Subsection I.4.F]. We assume a few standard properties from those references in the sequel. In particular, if $X$ is compact and metrizable, then so is $\mathcal{K}(X)$, for instance using a Hausdorff metric. As a result, in this case we can describe the Vietoris topology in terms of convergence of sequences. More specifically, for any sequence $(K_{n})_{n\geq 1}$ in $\mathcal{K}(X)$, define its topological upper and lower limits to be the sets
@@
respectively. If $X$ is compact and metrizable, then the sequence $(K_{n})_{n\geq 1}$ Vietoris converges if and only if its topological upper and lower limits are equal, and in this case that common set is the limit of the sequence [92, Subsection I.4.F].
We now endow the product space on the right-hand side of (8.4) with the product of the Vietoris topologies. Each $\Sigma_{k}(\mathfrak{A})$ is compact by the Banach–Alaoglu theorem and metrizable because $\mathfrak{A}$ is separable. Therefore each of those Vietoris topologies is again compact and metrizable, and hence so is that whole countable product topology.
@@
As far as I know, this topology was first studied by Abért and Elek in [1]. They were prompted by an analogous idea in graph theory from [20] (see also Remark 8.13 below). Abért and Elek focused on the analogous definition in ergodic theory, but their paper indicates the story for unitary representations of groups as well. However, the term ‘strong-quotient’ is new. We comment on this further after presenting some basic properties.
One key conclusion of the representation theoretic analog of [1, Theorem 1] is this: if one allows all separable representations, then the image of the summary map is a compact subset of the right-hand side of (8.4), and hence closed. The proof adapts with only cosmetic changes to a general separable unital C*-algebra $\mathfrak{A}$, and the result has the following consequence.
@@
Proof.
By the equivalence (8.1), if $\varphi\in\Sigma_{k}(\mathfrak{A})$, then it is asymptotically associated to $(\pi_{n})_{n\geq 1}$ if and only if it lies in $\mathrm{T}\limsup_{n}\overline{\Sigma_{k}(\pi_{n})}$. The choice of Tlim $\mathrm{T}\limsup$ rather Tlim $\mathrm{T}\liminf$ here is a matter of convention. If $(\pi_{n})_{n\geq 1}$ strong-quotient converges then this choice makes no difference, but otherwise it can be consequential. The next lemma illustrates this.
Lemma 8.12.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. Assume that $\boldsymbol{\pi}$ strong-quotient converges. If $\varphi$ and $\psi$ are disjoint and they are both asymptotically associated to $\boldsymbol{\pi}$, then so is $\mathrm{diag}(\varphi,\psi)$, and if $k=\ell$ then so is $\varphi+\psi$.
@@
If $k=\ell=1$, then the sum $\varphi+\psi$ is handled in the same way, using Corollary 8.8(a) in place of Lemma 8.6. Finally, the case when $k=\ell>1$ follows by Lemma 8.11. ∎
Without assuming strong-quotient convergence, the two factors in the Cartesian product above might be nonempty only along disjoint subsequences of indices $n$, and then the conclusion of the lemma may fail.
The term ‘strong-quotient’ is chosen to locate this topology relative to some other standard notions for representations. Let us quickly review some of these in order to explain this choice.
The most common topology on the spectrum $\widehat{\mathfrak{A}}$ is the Fell topology, which is the pullback of the Jacobson topology on $\mathrm{Prim}\,\mathfrak{A}$. This has several equivalent definitions, originally explored by Godemont [73] and Fell [59]: see [48, Chapter 3] or [62, Section 7.2] for textbook accounts.
One can try to extend this topology to larger spaces of representations, not just the irreducible ones. But then the different definitions no longer coincide in general. Two possibilities were defined and studied by Fell in [58, Section 3] and [60]. He called them the ‘quotient’ and ‘inner hull-kernel’ topologies. Modern practice usually assigns Fell’s name to his ‘inner hull-kernel topology’, but it is actually the quotient topology that is closer to our present needs.
These topologies can be defined by approximately comparing the sets of completely positive maps associated to representations. The difference boils down to whether one allows approximation by finite sums (giving the Fell topology) or not (giving the quotient topology). This sits alongside our discussion of the difference between weak and approximate association in Subsection 8.1.
Focusing on the quotient topology, it can be defined by pulling back a product topology on the space in (8.4) under the summary map, much like our approach to the strong-quotient topology. But the quotient topology uses the ‘lower Vietoris’ topology on each of the factors $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ (see [114, Appendix]), rather than the Vietoris topology. The lower Vietoris topologies are much weaker: indeed, they are usually not even T1, and so limits of sequences need not be unique. Concretely, a sequence $(\pi_{n})_{n\geq 1}$ converges to another representation $\pi$ in the quotient topology if
@@
So, in this topology, any representation whose $k$-summaries are all sufficiently small is a limit of $(\pi_{n})_{n\geq 1}$.
Another manifestation of the weakness of the quotient topology is the following. For a set $R$ of representations (or equivalence classes of them) and an element $a\in\mathfrak{A}$, the norm map
@@
is lower semicontinuous for the quotient topology on $R$, but not necessarily continuous. This is already true for $R=\widehat{\mathfrak{A}}$: see, for instance, the longer discussion in [59, Chapter II] or [48, Section 3.3].
To avoid that weakness, recent research has seen a swell of interest in ‘strong convergence’ of representations. This mode of convergence is defined precisely by requiring that the norm maps in (8.6) converge to their values on a limiting representation. See, for example, the survey [113].
@@
  • For quotient-topology convergence: for every open subset $O$ of $\Sigma_{k}(\mathfrak{A})$ that meets $\overline{\Sigma_{k}(\pi)}$, and for all sufficiently large $n$, there must be some $k$-tuples $V_{n}$ in $H_{\pi_{n}}$ such that $\Phi^{\pi_{n}}_{V_{n}}\in O$.
  • For strong convergence: for every $a\in\mathfrak{A}$, every $\varepsilon>0$, and all sufficiently large $n$, all unit vectors $v\in H_{\pi_{n}}$ must satisfy $\|\pi_{n}(a)v\|\leq\|\pi(a)\|+\varepsilon$.
On the other hand, control of the norm maps from (8.6) does not require controlling multiplicities, or tuples of more than one vector. In fact, the strong convergence of an AP sequence $(\pi_{n})_{n\geq 1}$ is equivalent to Vietoris convergence of the sequence of closed convex hulls $\overline{\mathrm{conv}}\,\Sigma_{1}(\pi_{n})$.
@@
8.4 Almost periodic sequences
In the next section, the almost periodic entropy of a completely positive map $\varphi$ is defined in terms of the volumes of the $O$-typical sets for certain representations as $O$ ranges over neighbourhoods of $\varphi$. The representations that we use belong to the following sequences.
@@ 8.4 Almost periodic sequences
Later we also need another mode of convergence for an AP sequence $\boldsymbol{\pi}$. This is the convergence of the tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ to some limit tracial state $\tau$ in the weak topology of $\mathfrak{A}_{+}^{\ast}$. In the terminology of free probability theory, this asserts that, for any finite subset $F$ of $\mathfrak{A}$, the tuples $(\pi_{n}(a):\ a\in F)$ form a sequence of ‘microstates’ for $F$ in the ‘non-commutative probability space’ $(\mathfrak{A},\tau)$ (see [156], for example). This convergence of tracial states is one of the hypotheses of Theorem C. It determines which limiting tracial state should be used to define the Fuglede–Kadison determinant in the conclusion of that theorem.
@@
After identifying $\overline{\mathrm{tr}}_{kd_{n}}$ with $\overline{\mathrm{tr}}_{d_{n}}\otimes\overline{\mathrm{tr}}_{k}$, this follows from the continuity of the operation $(\cdot)\otimes\overline{\mathrm{tr}}_{k}$ on tracial states. That continuity can be checked entry-wise against elements of $\mathrm{M}_{k}(\mathfrak{A})$. ∎
@@ 9 Almost periodic entropy
In this section we define the almost periodic entropy of an $\mathrm{M}_{k}$-valued completely positive map on a separable, unital C*-algebra $\mathfrak{A}$. We then build up its properties towards the proof of Theorem C, and follow that with a few consequences.
@@ 9.1 Preliminary results from high-dimensional probability
Then $v(d)$ is equal to the volume of the unit ball in $\mathbf{C}^{d}$ [127, Subsection 1.4.9]. By Stirling’s approximation, this function satisfies
@@
Recall that we write $\sigma_{2d-1}$ for the normalized spherical measure on the unit sphere $\mathrm{S}^{2d-1}$ in $\mathbf{C}^{d}$ or in any other space with a standing identification with $\mathbf{R}^{2d}$. A key feature of these measures in high dimensions is the phenomenon of measure concentration. We need the following special case.
@@
If $f(v)=\langle Av,v\rangle$ for $v\in\mathrm{S}^{2d-1}$, then $f$ is $2\|A\|$-Lipschitz, and its spherical average is $\overline{\mathrm{tr}}_{d}A$ by the invariance of trace under conjugation. Therefore the desired inequality follows from the concentration of general Lipschitz functions on high-dimensional spheres. See, for instance, [115, Corollary V.2] for a proof of this concentration result from its Gaussian analog, or [102] for a more modern proof by interpolation along the heat semigroup. ∎
@@
Lemma 9.14 below is a generalization of Corollary 9.2 that depends on Lemma 9.1 in the same way.
Subsection 1.2 discusses the analogy between the Gram matrix of a tuple of vectors and the joint distribution of a tuple of finite-valued random variables. The next theorem adds another layer to this analogy: a ‘method of types’ interpretation for $\log\det Q$ when $Q$ is a positive semi-definite matrix. It approximates the volumes of the sets $T(d,O)$ when $O$ is a small neighbourhood of $Q$. In information theory, the method of types captures the basic combinatorial meaning of discrete Shannon entropy [40, Section 11.1], and similarly the large-deviations interpretation of Kullback–Leibler divergence [41, Subsection 2.1]. The differential entropy of a jointly Gaussian random vector also has a method-of-types interpretation in terms of volumes [40, Section 8.2].
@@
    (Lower bound) If $O$ is any neighbourhood of $Q$ in $\mathrm{M}_{k+}$, then
@@
Variants of Theorem 9.3 are widely known, for instance in the study of complex Wishart distributions for random Gram matrices [74]. But I have not found a convenient reference for this particular one, so include a full proof.
@@
Identifying $\mathrm{M}_{d,k}$ with $\mathbf{C}^{dk}$ and integrating in polar coordinates [127, Subsection 1.4.3], we obtain
@@
By Corollary 9.2 and the asymptotic (9.2), this is greater than $v(d)^{k}$ for all sufficiently large $d$.
@@
If we regard a $d$-by-$k$ matrix as the $d$-tuple of its rows, then right-multiplication by $R$ on $\mathrm{M}_{d,k}$ becomes the direct sum of $d$ copies of $R^{\mathrm{T}}$ acting on $\mathbf{C}^{k}$. Regarded as a real linear transformation acting on $2kd$ real linear dimensions, this has Jacobian $|\det R|^{2d}$: see, for instance, [127, Subsection 1.3.5]. Therefore (9.5) gives
@@
Since $O^{\prime}$ is an arbitrary neighbourhood of $Q^{\prime}$, and
$\det Q^{\prime}=|\det R|^{2}\cdot\det Q,$
this shows that part (a) for $Q^{\prime}$ follows from part (a) for $Q$. Similarly, if $R$ is invertible, then we may reverse the roles of $Q$ and $Q^{\prime}$ above and deduce that part (b) for $Q^{\prime}$ follows from part (b) for $Q$.
@@
$O:=\{Q^{\prime}\in\mathrm{M}_{k+}:\ q_{11}^{\prime}<\varepsilon^{2}\ \hbox{and}\ \max_{i}q^{\prime}_{ii}<r^{2}\}.$
Then $O$ is a neighbourhood of $Q$, and
$T(d,O)=\big{\{}X\in\mathrm{M}_{d,k}:\ \|x_{1}\|<\varepsilon\ \hbox{and}\ \max_{i}\|x_{i}\|<r\big{\}}.$
Therefore this neighbourhood satisfies
@@
Theorem 9.3 is a template for Theorem C (and also a special case of that theorem with $\mathfrak{A}=\mathrm{M}_{k}$). The rather soft proof given above, which avoids evaluating any integrals over spaces of matrices exactly, is also a precursor to the proof of Theorem C.
For discrete Shannon entropy, the usual proofs in the method of types involve counting strings that have exactly a given empirical distribution and then applying Stirling’s approximation. By contrast, in Step 3 of the proof of Theorem 9.3, we use a change of variables to transport the desired estimate from $I_{k}$ to most other elements of $\mathrm{M}_{k+}$. This way of using the symmetries of $\mathrm{vol}_{2kn}$ has no obvious analog for probability distributions over finite sets. A related use of symmetry is also essential to our later proof of Theorem C.